### v1

@@ I Introduction
Labyrinthine patterns, characterized by intricate stripe-like features, are hallmarks of non-linear systems that are ubiquitous in nature and appear in biological patterns, chemical reactions, crystal growth, and magnetic ordering [1, 2, 3, 4, 5]. Fig. 1 illustrates a labyrinthine pattern in a magnetic film. These patterns consist of well-defined local stripes whose orientations vary randomly over large regions and are disrupted by numerous defects, resulting in the absence of long-range order. Because of this lack of long-range correlation, defining a conventional order parameter to characterize their structure is challenging, so new methods of characterization beyond traditional phase descriptions are required.
Indeed, numerous attempts have been made to characterize labyrinthine patterns by using, e.g., structure factors (Fourier transform of labyrinthine patterns) [6, 7, 8, 9, 10] and disorder functions [11, 12, 13, 14]. These methods utilize the global features of labyrinthine patterns and have provided simple measures of disordered structures. Meanwhile, they can potentially fail to capture the local structural properties such as the defects in local stripes despite their importance in disordered systems.
This difficulty hinders the understanding of the physical properties associated with such patterns, including magnetic, electric, and thermal properties, and thus a new and efficient characterization method applicable to such a disordered system is highly anticipated in material science, prompting numerous attempts thus far [7, 6, 8, 9, 11, 12, 14, 15, 16, 17, 18].
In stripe pattern formation, defects generally exhibit unique geometrical properties associated with symmetry breaking of the local order parameter, and these properties impose strong geometrical constraints on the evolution of patterns. In the case of labyrinthine patterns, two different types of defects called junctions and terminals are known to play crucial roles in their formation [16, 17, 18]. A single defect cannot be removed by a continuous deformation of labyrinthine patterns, while pairs of different defects can be pair annihilated due to their geometrical properties. Typically, physical systems evolve over time toward energetically favorable, less disordered configurations. However, due to the necessity of pair annihilation of defects for transitions to more ordered states, global structural reconfigurations become kinetically suppressed. This leads to stationary labyrinthine patterns, despite the presence of numerous defects and the absence of long-range order.
@@
To characterize the junctions and terminals in Bi:YIG films, Okubo et al. [19] developed TM-CNN, a technique that leverages template matching and CNN (Convolutional Neural Network) classification for accurate topological defect identification. Building on this technique, Shimizu et al. [20] conducted a systematic analysis of the topological defects, demonstrating that their spatial correlations serve as an effective metric for capturing the spatial evolution of labyrinthine patterns.
@@
We combine image restoration (IR) and segmentation techniques to train a model on synthetically degraded images, achieving accurate segmentation of the stripe pattern formed on Bi:YIG films. While degradations such as blur, rain, and Gaussian noise are commonly used within image restoration, we demonstrate that Simplex noise can simulate smudge-like occlusions. The resulting segmentations were mapped into graph representations of the magnetic pattern, efficiently storing topological defect locations along with the length and curvature of the stripe segments connecting them. Through geometric analysis, we characterize the propagation of magnetic stripes, uncovering new aspects of their geometric and topological organization. In particular, we characterize an interleaved behavior resulting from the field annealing protocol, find particular geometric properties in the quenched to annealed transition, and identify geometric differences between junction-junction and junction-terminal stripe segments.
@@ II Related works
Image segmentation is essentially classifying each pixel in an image. Long et al. [21] proposed Fully Convolutional Networks, a fully convolutional architecture with end-to-end training for image segmentation. It employs an encoder-decoder architecture, sequentially generating higher-level features from the input image until applying segmentation. The U-Net model proposed by Ronneberger et al. [22] introduced a U-shaped encoder-decoder CNN architecture for image segmentation. Through the usage of concatenation bridges, the network can leverage jointly detailed spatial information from low-level features and semantic context from high-level features. This structure enables highly sample-efficient training while maintaining a low memory footprint relative to the number of parameters.
@@
Since its introduction, U-Net has become a widely used architecture across various tasks. It is particularly popular for medical image segmentation [23] and has also demonstrated success in general microscopy image segmentation [24]. Still, most research on image segmentation has focused on clean and well-defined images. Some studies have addressed various types of degradations, such as white noise, blurring, haze, rain, and low-light conditions [25, 26, 27]. However, image segmentation in settings with visual obstructions or occlusions remains relatively underexplored.
@@
Degradations are more commonly tackled in the field of image restoration, which is dedicated in obtaining high-quality, clear images from degraded ones. Most deep learning approaches rely on synthetic degradations to create clean and degraded image pairs for model training. These degradations do not necessarily reflect real-world conditions, so creating more representative degradations remains a challenge, motivating the development of many domain-specific approaches [28, 29, 30, 31].
@@
Additive White Gaussian Noise (AWGN) filtering is a major area of interest in image restoration. Early neural network-based approaches, such as Multi-Layer Perceptrons (MLPs), showed promise in noise filtering tasks by matching the performance of previous techniques at a single noise level, yet they struggled with variable noise levels [32]. Since then, CNNs have not only become widely adopted for AWGN filtering, but have also proved highly successful. Zhang et al. [33] introduced DnCNN, a CNN model designed for Gaussian noise removal through residual image extraction. Later, Zhang et al. [34] proposed FFDNet, which further improved Gaussian noise filtering by incorporating noise level maps as inputs, enabling better adaptability to varying noise levels across an image. Park et al. [35] introduced a U-Net-based model for noise filtering. To handle different noise levels, their model was trained using Gaussian noise augmentation with varying variance settings.
@@
In contrast, Simplex and Simplex-like noises have been used more sporadically in task-specific settings. Perlin noise, which is similar to simplex noise, was proposed for generating random Region of Interest patches in HRCT images [36]. In chest X-ray classification of lung cancer, Perlin noise has also been combined with AWGN as an augmentation strategy [37]. Beyond medical imaging, Perlin noise has been adopted to simulate haze for dehazing tasks [38]. Mahmood et al. [39] applied simplex noise to define the shape of synthetic cracks for segmentation of structural cracks. More recently, it has been employed to generate anomalies for industrial image anomaly detection [40].
@@
Transformer-based models have recently demonstrated superior generalization capabilities compared to CNNs when trained with sufficient data and computational resources [41]. Consequently, many modern computer vision models now leverage transformer-based architectures, in both segmentation [42, 43, 44] and IR [45, 46, 47]. Alternative techniques for IR have also been proposed by leveraging generative models. Isola et al.[48] introduced pix2pix, a framework based on Generative Adversarial Networks (GANs) [49] for general image-to-image translation. Their flexible approach allows for its usage in both image restoration and image segmentation tasks. Alternatively, Majurski et al. used trained GAN discriminators for pre-optimizing U-Nets for segmenting cell images[50]. Diffusion models have also been studied for image restoration[51, 52, 53].
@@
Deocclusion is a subfield within image restoration commonly associated with inpainting. Many single-image deocclusion approaches rely on a two-step procedure: first identifying occlusion locations, then inpainting the affected areas. Several methods follow this two-stage approach. For example, Matsui and Ikehara [29] utilized a U-Net model to segment fence-occluded regions. For raindrop removal, Qian et al. [54] introduced a GAN-based method that generates a heatmap of potential raindrop locations and subsequently produces an image without raindrops using conditional generation. Dong et al.[55] also adopted a two-step framework for face deocclusion, combining U-Nets with GAN training for occlusion segmentation and inpainting. Other approaches perform deocclusion in an unified model. Angah and Chen [56] approached construction site deocclusion by directly training a single U-Net model using GAN framework for training. In the case of face deocclusion, Zhao et al. [28] proposed a model that uses an LSTM-Autoencoder trained on paired clean and synthetically occluded images, jointly learning to detect and restore occlusions.
@@ III Methodology
To analyze the magnetic labyrinthine pattern, we developed a processing pipeline illustrated in Fig. 3. Segmentation of the light and dark regions is performed through a U-Net model trained to filter both noise and occlusion degradations. Two distinct types of paths are then characterized:
@@ III-A Dataset
Images of labyrinthine magnetic domains were acquired from thin films of bismuth‑doped yttrium iron garnet subjected to a controlled magnetic‑field annealing protocol. Each annealing trial began by saturating the film’s magnetization with a strong external out‑of‑plane field ($\pm z$). The field was then switched off and held at zero for $10$ s, allowing the magnetization to relax into a domain structure with opposing polarities. The resulting domain pattern was captured using polarized‑light microscopy[57].
@@
Each sequence type was repeated six times, yielding 12 complete trials and a total of 444 domain images with $5200\times 3888$ resolution. All measurements were performed at room temperature on a 2 mm $\times$ 1.8 mm film.
@@ III-B Segmentation
A U-Net-based model was used to segment the magnetic stripe pattern images. Initially, we attempted a simple local Otsu thresholding algorithm [58]. While it properly segmented clear regions of the magnetic stripe pattern, the results were unsatisfactory in areas affected by noise or occlusion degradation. Table I shows a side-by-side comparison of the Otsu and U-Net results on both clear and degraded patches.
TABLE I: Comparison of segmentation methods across clear, occluded, and noisy regions. U-Net outperforms Otsu thresholding, especially under noise and occluded conditions.
@@
Noisy [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ III-B1 Pre-processing
We then extracted $320\times 320$ patches from low-noise regions, where light and dark areas are clearly distinct. In these well-defined regions, local Otsu thresholding produced accurate segmentations that served as targets for the U-Net. To locate the low-noise areas systematically, we computed local brightness variance with a sliding window of $64\times 64$ pixels, advancing in 32-pixel steps. Regions with pronounced bright and dark contrast exhibited higher variance, making them easy to separate from noisy zones.
The resulting low-resolution variance map was upscaled to the original $5200\times 3888$ resolution with bilinear interpolation. We sampled patches only where the local variance was above $80\%$ of the image-wide variance, ensuring that most selected areas were well defined, though minor degradations remained.
Fig. 5 shows the outcome: yellow marks the low-noise regions used for training, while purple highlights the high-noise areas that were excluded.
@@
In order to make the model capable at segmenting the degraded regions of the magnetic stripe pattern, we synthetically introduced degradations during training to the labyrinthine patches using Gaussian and Simplex noises [59]. Simplex noise allowed for the creation of smooth regions with correlated values, which was used to simulate occlusion degradation. To represent the diverse noise intensities and structural distortions found in the original images, the noise parameters were randomly sampled as an augmentation step during training (Fig. 6). This ensured that the network learned to filter images with different noise characteristics.
Appendix VI-A presents additional details on the synthetic degradations and includes a comparative table illustrating their impact on segmentation accuracy.
Figure 6: Examples of a magnetic stripe patch with different synthetic noise configurations added.

### v2

@@ I Introduction
Labyrinthine patterns, characterized by intricate stripe-like features, are hallmarks of non-linear systems that are ubiquitous in nature and appear in biological patterns, chemical reactions, crystal growth, and magnetic ordering [Nicolis77, Cross93, Koch94, Pismen06, Cross09]. Fig. 1 illustrates a labyrinthine pattern in a magnetic film. These patterns consist of well-defined local stripes whose orientations vary randomly over large regions and are disrupted by numerous defects, resulting in the absence of long-range order. Because of this lack of long-range correlation, defining a conventional order parameter to characterize their structure is challenging, so new methods of characterization beyond traditional phase descriptions are required.
Indeed, numerous attempts have been made to characterize labyrinthine patterns by using, e.g., structure factors (Fourier transform of labyrinthine patterns) [Ouyang91, Elder92, Cross95, Christensen98, Alar20] and disorder functions [Gunaratne95, Hu04, Nathan05, Hu05]. These methods utilize the global features of labyrinthine patterns and have provided simple measures of disordered structures. Meanwhile, they can potentially fail to capture the local structural properties such as the defects in local stripes despite their importance in disordered systems.
This difficulty hinders the understanding of the physical properties associated with such patterns, including magnetic, electric, and thermal properties, and thus a new and efficient characterization method applicable to such a disordered system is highly anticipated in material science, prompting numerous attempts thus far [Elder92, Ouyang91, Cross95, Christensen98, Gunaratne95, Hu04, Hu05, Hou97, Seul92, Seul92b, Seul92c].
In stripe pattern formation, defects generally exhibit unique geometrical properties associated with symmetry breaking of the local order parameter, and these properties impose strong geometrical constraints on the evolution of patterns. In the case of labyrinthine patterns, two different types of defects called junctions and terminals are known to play crucial roles in their formation [Seul92, Seul92b, Seul92c]. A single defect cannot be removed by a continuous deformation of labyrinthine patterns, while pairs of different defects can be pair annihilated due to their geometrical properties. Typically, physical systems evolve over time toward energetically favorable, less disordered configurations. However, due to the necessity of pair annihilation of defects for transitions to more ordered states, global structural reconfigurations become kinetically suppressed. This leads to stationary labyrinthine patterns, despite the presence of numerous defects and the absence of long-range order.
@@
To characterize the junctions and terminals in Bi:YIG films, Okubo et al. [okubo2024] developed TM-CNN, a technique that leverages template matching and CNN (Convolutional Neural Network) classification for accurate topological defect identification. Building on this technique, Shimizu et al. [shimizu2023] conducted a systematic analysis of the topological defects, demonstrating that their spatial correlations serve as an effective metric for capturing the spatial evolution of labyrinthine patterns.
@@
We combine image restoration (IR) and segmentation techniques to train a model on synthetically degraded images, achieving accurate segmentation of the stripe pattern formed on Bi:YIG films. While degradations such as blur, rain, and Gaussian noise are commonly used within image restoration, we demonstrate that Simplex noise [perlin2002] can simulate smudge-like occlusions. The resulting segmentations were mapped into graph representations of the magnetic pattern, efficiently storing topological defect locations along with the length and curvature of the stripe segments connecting them. Through geometric analysis, we characterize the propagation of magnetic stripes, uncovering new aspects of their geometric and topological organization. In particular, we characterize an interleaved behavior resulting from the field annealing protocol, find particular geometric properties in the quenched to annealed transition, and identify geometric differences between junction-junction and junction-terminal stripe segments.
@@ II Related works
Image segmentation is essentially classifying each pixel in an image. Long et al. [long2015] proposed Fully Convolutional Networks, a fully convolutional architecture with end-to-end training for image segmentation. It employs an encoder-decoder architecture, sequentially generating higher-level features from the input image until applying segmentation. The U-Net model proposed by Ronneberger et al. [ronneberger2015] introduced a U-shaped encoder-decoder CNN architecture for image segmentation. Through the usage of concatenation bridges, the network can leverage jointly detailed spatial information from low-level features and semantic context from high-level features. This structure enables highly sample-efficient training while maintaining a low memory footprint relative to the number of parameters.
@@
Since its introduction, U-Net has become a widely used architecture across various tasks. It is particularly popular for medical image segmentation [azad2024] and has also demonstrated success in general microscopy image segmentation [Wu2022]. Still, most research on image segmentation has focused on clean and well-defined images. Some studies have addressed various types of degradations, such as white noise, blurring, haze, rain, and low-light conditions [xia2019, guo2019, rajagopalan2023]. However, image segmentation in settings with visual obstructions or occlusions remains relatively underexplored.
@@
Degradations are more commonly tackled in the field of image restoration, which is dedicated in obtaining high-quality, clear images from degraded ones. Most deep learning approaches rely on synthetic degradations to create clean and degraded image pairs for model training. These degradations do not necessarily reflect real-world conditions, so creating more representative degradations remains a challenge, motivating the development of many domain-specific approaches [zhao2017, matsui2019, sakaridis2018, hao2019].
@@
Additive White Gaussian Noise (AWGN) filtering is a major area of interest in image restoration. Early neural network-based approaches, such as Multi-Layer Perceptrons (MLPs), showed promise in noise filtering tasks by matching the performance of previous techniques at a single noise level, yet they struggled with variable noise levels [burger2012]. Since then, CNNs have not only become widely adopted for AWGN filtering, but have also proved highly successful. Zhang et al. [zhang2017] introduced DnCNN, a CNN model designed for Gaussian noise removal through residual image extraction. Later, Zhang et al. [zhang2018] proposed FFDNet, which further improved Gaussian noise filtering by incorporating noise level maps as inputs, enabling better adaptability to varying noise levels across an image. Park et al. [park2019] introduced a U-Net-based model for noise filtering. To handle different noise levels, their model was trained using Gaussian noise augmentation with varying variance settings.
@@
In contrast, Simplex and Simplex-like noises have been used more sporadically in task-specific settings. Perlin noise [perlin1985], which is similar to simplex noise, was proposed for generating random Region of Interest patches in HRCT images [Bae2018]. In chest X-ray classification of lung cancer, Perlin noise has also been combined with AWGN as an augmentation strategy [Haekal2021]. Beyond medical imaging, Perlin noise has been adopted to simulate haze for dehazing tasks [Zheng2024]. Mahmood et al. [Mahmood2024] applied simplex noise to define the shape of synthetic cracks for segmentation of structural cracks. More recently, it has been employed to generate anomalies for industrial image anomaly detection [Li2025].
Transformer-based models demonstrate superior generalization capabilities compared to CNNs when trained with sufficient data and computational resources. Transformers allow global interactions between input tokens through the attention mechanism, while convolution-based approaches are biased toward local patterns due to their translation-invariant receptive fields. Vision Transformers (ViT) were introduced by Dosovitskiy et al. [dosovitskiy2020] for image classification, representing images as sequences of patch embeddings processed by a pure transformer encoder. Building on ViT, several transformer-based architectures have been proposed for segmentation. SETR [zheng2021] uses a ViT encoder with convolutional decoders for semantic segmentation, while TransUNet [chen2021Transunet] combines a CNN encoder with transformer blocks and a U-Net-style decoder. Segmenter [strudel2021] employs a ViT backbone and a transformer-based decoder to produce class masks. Swin Transformer [liu2021] introduces a hierarchical windowed attention backbone, which has been adapted to a U-shaped design in Swin-Unet [cao2022] for segmentation. SegFormer [xie2021segformer] is a hierarchical transformer-based model whose encoder produces multiscale features using overlapping patch embeddings, while a simple all-MLP decoder upsamples and fuses features into high-resolution segmentation maps without positional encodings.
@@
Alternative techniques for IR have also been proposed by leveraging the transformer architecture  [chen2021, liang2021, zamir2022] and generative models. Isola et al.[isola2017] introduced pix2pix, a framework based on Generative Adversarial Networks (GANs) [goodfellow2014] for general image-to-image translation. Their flexible approach allows for its usage in both image restoration and image segmentation tasks. Alternatively, Majurski et al. used trained GAN discriminators for pre-optimizing U-Nets for segmenting cell images[majurski2019]. Diffusion models have also been studied for image restoration[kawar2022, luo2023, xie2023].
@@
Deocclusion is a subfield within image restoration commonly associated with inpainting. Many single-image deocclusion approaches rely on a two-step procedure: first identifying occlusion locations, then inpainting the affected areas. For example, Matsui and Ikehara [matsui2019] utilized a U-Net model to segment fence-occluded regions. For raindrop removal, Qian et al. [qian2018] introduced a GAN-based method that generates a heatmap of potential raindrop locations and subsequently produces an image without raindrops using conditional generation. Dong et al.[dong2020] also adopted a two-step framework for face deocclusion, combining U-Nets with GAN training for occlusion segmentation and inpainting. Other approaches perform deocclusion in an unified model. Angah and Chen [angah2020] approached construction site deocclusion by directly training a single U-Net model using GAN framework for training. In the case of face deocclusion, Zhao et al. [zhao2017] proposed a model that uses an LSTM-Autoencoder trained on paired clean and synthetically occluded images, jointly learning to detect and restore occlusions.
@@ III Methodology
To analyze the magnetic labyrinthine pattern, we developed a processing pipeline illustrated in Fig. 3. Segmentation of the light and dark regions is performed through a U-Net model.
The U-Net model was chosen for its widespread use in microscopy segmentation [Wu2022]. Although Transformer-based approaches have recently gained popularity due to their ability to model global context, the propagation of labyrinthine patterns is primarily governed by their local structure. Similarly, noise and occlusion degradations in labyrinthine images are spatially confined. U-Net’s local inductive bias is thus a good match for this task, supporting strong generalization in this data-constrained setting.
Two distinct types of paths are then characterized:
@@ III-A Dataset
Images of labyrinthine magnetic domains were acquired from thin films of bismuth‑doped yttrium iron garnet subjected to a controlled magnetic‑field annealing protocol. Each annealing trial began by saturating the film’s magnetization with a strong external out‑of‑plane field ($\pm z$). The field was then switched off and held at zero for $10$ s, allowing the magnetization to relax into a domain structure with opposing polarities. The resulting domain pattern was captured using polarized‑light microscopy[Puchalska78].
@@
Each sequence type was repeated six times, yielding 12 complete trials and a total of 444 domain images with $5200{\times}3888$ resolution. Images can then be indexed by the tuple:
$I\equiv(t,\sigma,s)$ (1)
$t\in\{1,\dots,6\}$ denotes the trial, $\sigma\in\{+,-\}$ is the sequence type (positive/negative), and $s\in\{0,\dots,36\}$ is the magnetic field annealing step index. All measurements were performed at room temperature on a $2\text{mm}{\times}1.8\text{mm}$ film. Table I summarizes the dataset characteristics.
TABLE I: Dataset overview.
Index Parametrization
Trial index ($t$) $\{1,2,3,4,5,6\}$
Sequence ($\sigma$) $\{+,-\}$
Step index ($s$) $\{0,1,2,\dots,36\}$
Types in Positive Sequence ($\sigma\in\{+\}$), indexes $s$
Type A $\{0,2,4,\dots,36\}$
Type B $\{1,3,5,\dots,35\}$
Types in Negative Sequence ($\sigma\in\{-\})$, indexes $s$
Type A $\{1,3,5,\dots,35\}$
Type B $\{0,2,4,\dots,36\}$
Image properties
Resolution $5200\text{px}{\times}3888\text{px}$
Film size $2\text{mm}{\times}1.8\text{mm}$
Total images $(6+6)\times 37=444$
@@ III-B Segmentation
Prior to U-Net-based segmentation, we attempted a simple local Otsu thresholding algorithm, which determines local thresholds by maximizing the inter-class variance between intensity domains [otsu1975]. Table II shows a side-by-side comparison of the Otsu and U-Net segmentations on both clear and degraded patches.
In magnetic stripe patterns, regions with high local intensity variance exhibit a greater intensity separation between domains and a clearer bimodal intensity distribution (Table II, first row). As a result, segmentation in such regions is generally consistent with the visually identifiable stripe morphology.
However, in regions with reduced illumination, brightness and contrast decrease, making the intensity distributions of magnetic domains less separable (Table II, second row). This reduced contrast amplifies the relative impact of sensor noise, hindering threshold-based segmentation. Additionally, structured occlusions partially mask the stripe pattern, introducing localized distortions in the segmentation. Although occlusions may appear throughout the image, in practice they tend to introduce textures with intermediate intensities, reducing the local variance as well (Table II, third row).
The U-Net model was then developed to perform robust segmentation under these degradations. Training was performed using high-variance regions of the image with stronger domain separability, and the corresponding Otsu segmentations were used as pseudo-labels (“ground-truth proxies”). Synthetic degradations were then applied to the input images while keeping the pseudo-labels fixed, so that the network could learn to recover the stripe structure under noise and occlusion without inheriting the limitations of threshold-based segmentation.
TABLE II: Comparison of segmentation methods across clear, noisy, and occluded regions. U-Net outperforms Otsu thresholding in settings under noise and occlusions.
@@
Noisy [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ III-B1 Pre-processing
We then extracted $320{\times}320$ patches from the high-contrast regions. To locate these areas systematically, we computed local brightness variance with a sliding window of $64{\times}64$ pixels, advancing in 32-pixel steps. The resulting low-resolution variance map was upscaled to the original $5200{\times}3888$ resolution with bilinear interpolation. We sampled patches only where the local variance was above $80\%$ of the image-wide variance. Fig. 5 shows the outcome: yellow marks the low-noise regions used for training, while purple highlights the high-noise areas that were excluded.
@@
In order to make the model capable of segmenting the degraded regions of the magnetic stripe pattern, we synthetically introduced degradations to the clear patches used for the training. To account for uneven illumination and contrast variations in the acquired images, we included photometric transformation augmentation consisting of brightness and gamma perturbations. AWGN was then used to model noise degradation, as a simple representation of uncorrelated sensor noise commonly assumed in imaging systems [gonzalez2018].
In contrast, occlusions must be smooth, spatially correlated, and textured to resemble real smudge-like artifacts. To generate such patterns procedurally, Perlin noise, widely used in texture generation within computer-generated imagery [ebert2003], provides a means to produce coherent, natural-appearing textures [perlin1985]. In practice, we employed Simplex noise [perlin2002], a derivative of the Perlin algorithm, to simulate occlusions due to its lower computational cost and reduced directionality artifacts.
To represent the diverse noise intensities and structural distortions found in the original images, the noise parameters were randomly sampled as an augmentation step during training (Fig. 6). This ensured that the network learned to segment images with different noise characteristics.
Appendix A presents additional details on the synthetic degradations and includes a comparative table illustrating their impact on segmentation accuracy.
Figure 6: (a) Original magnetic stripe patch. (b–f) Instances of the same patch after adding various configurations of synthetic noise.